OpenStudy (anonymous):

the area of the region bounded by lines x=0, x=2 and y=0 and the curve y=e^(x/2) is.... a)(e-1)/2 b)e-1 c)2(e-1) d)2e-1 e)2e please show me how to do this....

OpenStudy (anonymous):

integrate y(x) on the interval 0-2

OpenStudy (anonymous):

i know i need to do that but i dont know how to.

OpenStudy (turingtest):

a u-sub is needed\[u=\frac x2\implies du=\frac{dx}2\]

OpenStudy (anonymous):

where are you stuck?

OpenStudy (turingtest):

rewrite your integral in terms of u and let us know where you get stuck

OpenStudy (anonymous):

the derivative of e^x/2 is what confused me

OpenStudy (anonymous):

d(e^u)/du=e^u

OpenStudy (anonymous):

so when i write the integral it would be from 0 to 1 (e^u)du

OpenStudy (anonymous):

oh with a 2 in front though?

OpenStudy (turingtest):

if you want to evaluate it in terms of u, then yes

OpenStudy (anonymous):

then i have to do what?

OpenStudy (turingtest):

write down the answer... there's nothing to do after you evaluate, that's the last step

OpenStudy (anonymous):

i thought i have to integrate to get rid of the du?

OpenStudy (anonymous):

yes, substitution, integration, then evaluation. you can either evaluate for u if you change intervals or plug x"s back in for u and then evaluate

OpenStudy (anonymous):

\[2\int\limits_{0}^{1}e ^{u}du\]

OpenStudy (anonymous):

thats what i have

OpenStudy (anonymous):

now integrate

OpenStudy (anonymous):

sooo id get... umm 2* idk.

OpenStudy (anonymous):

i dont understand e's

OpenStudy (anonymous):

the integral of e^u*du is e^u

OpenStudy (anonymous):

evaluate for upper and lower limits of u or x

OpenStudy (anonymous):

ohh okay, thanks

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

i got 2(e-1)

OpenStudy (anonymous):

k